# OccFusion: A Straightforward and Effective Multi-Sensor Fusion Framework for 3D Occupancy Prediction

### 13 Mar 2024 v2

@@ III-C Surround-View Images Feature Extraction
Given surround-view images, we initially employ ResNet101-DCN [54] as our 2D backbone and feature pyramid network (FPN) [55] as the neck to extract multi-scale feature maps. The resulting feature maps have resolutions that are $\frac{1}{8}$, $\frac{1}{16}$, and $\frac{1}{32}$ of the input image resolution, respectively. Subsequently, the method proposed in [14] is utilized for view transformation, yielding multi-scale global BEV features and local 3D feature volumes. The global and local features with smaller resolutions contain valuable semantic information, aiding the model in predicting the semantic class of each voxel grid. Conversely, those features with larger resolutions provide rich spatial information, enabling the model to determine whether the current voxel grid is occupied or unoccupied.
@@ III-D Lidar Dense 3D Point Cloud Feature Extraction
In this paper, we adopt VoxelNet [32] as our 3D backbone for feature extraction of the 3D point cloud. We begin by voxelizing the 3D point cloud to generate the voxel grid and its associated coordinates. For each voxel grid that contains 3D points, we randomly select 35 points. Each point, denoted as $p_{i}^{Lid}$, includes an initial feature vector $p_{i}^{Lid}=[x_{i},y_{i},z_{i},\gamma_{i}]$ that represents the point’s 3D position and reflectance rate. Subsequently, we calculate the centre position of these 3D points and augment each point with a relative offset with respect to the centre position. This augmentation results in a new feature vector $p_{i}^{Lid}=[x_{i},y_{i},z_{i},\gamma_{i},x_{i}-\bar{x},y_{i}-\bar{y},z_{i}-\bar{z}]$. Following this, only the voxel grid with 3D points is input to the 3D backbone to refine the features further, producing the final local 3D feature volume. Additionally, for the global BEV feature, we employ average pooling along the Z-axis of the local 3D feature volume to obtain the flattened global BEV feature. The 3D backbone outputs the highest resolution of the global and local features, while the lower resolution features are obtained through 3D/2D downsampling operations.
@@ III-E Radar Sparse 3D Point Cloud Feature Extraction
The Radar 3D point contains richer information compared to the Lidar 3D point. Each Radar 3D point has an initial feature vector $p_{i}^{Rad}=[x_{i},y_{i},z_{i},V_{xi},V_{yi}]$, where $V_{xi}$ denote the velocity along the X-axis, and $V_{yi}$ denote the velocity along the Y-axis. Similar to the way used to process the Lidar point cloud, we begin with voxelizing the Radar 3D point cloud and obtain the voxel grids and their associate coordinates. For the voxel grid with points, we calculate the mean value among the 3D points and augment each point with a relative offset with respect to the mean value. This augmentation results in a new feature vector $p_{i}^{Rad}=[x_{i},y_{i},z_{i},V_{xi},V_{yi},x_{i}-\bar{x},y_{i}-\bar{y},z_{i}-\bar{z},V_{xi}-\bar{V_{x}},V_{yi}-\bar{V_{y}}]$. We subsequently input the non-empty voxel grid into the 3D backbone to obtain the local 3D feature volume and then apply average pooling to obtain the global BEV feature.
@@ III-F Dynamic Fusion 3D/2D
Drawing inspiration from BEVFusion [42, 43] and SENet [56], this research merges two BEV features and two 3D feature volumes by concatenating their feature channels. Subsequently, a Conv3D/2D layer is applied to reduce the feature channel dimension, facilitating the merging of valuable features from diverse modalities while filtering out noisy features. This process is followed by a 3D/2D SENet block, where the merged features are inputted into the squeeze module (SENet Block 3D/2D Upper) to determine the importance of each feature channel. The excitation module (SENet Block 3D/2D bottom) then implements an excitation procedure by multiplying the merged features with the squeeze feature, enabling critical features to dominate. Details of the Dynamic Fusion 3D module can be found in the upper section of Figure 3, while specifics of the Dynamic Fusion 2D module are depicted in the bottom section of Figure 3.
@@ IV-A Implementation Details
The OccFusion utilizes ResNet101-DCN [57, 58] as the 2D backbone, with pre-trained weights provided by FCOS3D [59], to extract image features. The feature maps from stage 1, 2, and 3 of the backbone are then fed into FPN [55], resulting in three levels of multi-scale image features. The network architecture consists of four levels ($L=4$), with no skip connection applied to the highest level. To extract 3D feature volumes, the VoxelNet [32] is used as the framework’s 3D backbone, which processes the lidar dense 3D point cloud and the sparse 3D point cloud from surround-view radars. For optimization, the AdamW optimizer with an initial learning rate of 5e-5 and weight decay of 0.01 is employed. The learning rate is decayed using a multi-step scheduler. During the training phase, the framework is trained using focal loss [60], Lovasz-softmax loss [61], and scene-class affinity loss [1]. To account for the importance of high-resolution 3D volumes over lower-resolution ones, a decayed loss weight $w=\frac{1}{2^{l}}$ is used for supervision at the $l$-th level. The model is trained on eight A10 GPUs, each with 24GB of memory, for six epochs to perform 3D semantic occupancy prediction task.
IV-B Dataset
@@
We conducted our experiments on the nuScenes dataset. The ground truth labels for 3D semantic occupancy prediction were obtained from SurroundOcc’s work [7], which offers a more comprehensive coverage than the well-known Occ3D [6] ground truth labels. The SurroundOcc’s label spans a range of -50m to 50m for the X and Y directions and -5m to 3m for the Z direction. This range is suitable for conducting our model’s ablation study on the perception range factor. As the test set labels were unavailable, we trained our model on the training set and evaluated its performance on the validation set. Furthermore, we selected specific frames from the validation set to create subsets for rainy and night scenarios, respectively. The distribution of classes in the validation set, rainy scenario subset, and night scenario subset is shown in Figure 4. In each set, the class with a sample number equal to zero is not listed.
IV-C Performance Evaluate Metrics
To evaluate the performance of different state-of-the-art (SOTA) algorithms and compare them with our approach in the task of 3D semantic occupancy prediction, we employ the intersection over union (IoU) to evaluate each semantic class. Additionally, we use the mean IoU across all semantic classes (mIoU) as a comprehensive evaluation metric:
@@ IV-C Performance Evaluate Metrics
IV-D Model Performance Analysis
We assess different sensor fusion strategies using our proposed framework on the nuScenes validation set, and present the final benchmark results in Table I. In comparison to purely vision-centric approaches, the inclusion of sparse 3D point cloud information from the surround-view radars leads to a noteworthy 3% improvement in performance. Moreover, the addition of dense 3D point cloud information from the lidar further enhances the performance to approximately 34% mIoU. These experimental results confirm the efficacy of utilizing multi-sensor fusion to substantially improve the performance of the 3D semantic occupancy prediction task.
Method Backbone Input Modality Resolution mIoU $\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ const. veh. $\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck $\bullet$ drive. surf. $\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation InverseMatrixVT3D [14] R101-DCN C $200\times 200\times 16$ 25.95 26.13 15.19 42.65 43.92 11.42 20.82 22.68 13.45 22.33 36.04 47.47 16.26 21.47 25.33 16.37 33.73 OccFusion R101+VoxelNet C+R $200\times 200\times 16$ 28.07 27.82 17.71 47.41 46.60 13.73 23.02 22.72 17.77 26.29 37.58 47.56 19.45 27.03 23.88 22.90 27.56 OccFusion R101+VoxelNet C+L $200\times 200\times 16$ 33.92 31.79 18.91 49.36 48.95 24.77 24.10 29.70 22.26 29.91 36.09 47.98 24.18 28.18 31.87 40.82 53.94 OccFusion R101+VoxelNet C+L+R $200\times 200\times 16$ 34.77 33.05 22.46 46.62 52.51 22.73 26.96 30.42 20.74 25.42 41.24 53.47 19.83 31.51 35.56 39.25 54.57
@@ IV-D Model Performance Analysis
IV-E Challenging Scenarios Performance Analysis
To gain deeper insights into sensor fusion properties and effectiveness, we assess various sensor fusion strategies in challenging nighttime and rainy scenarios. The performance of the models in these scenarios is presented in Table II and Table III.
In the rainy scenario, despite the radar sensor being the most resilient to adverse weather conditions, integrating surround-view camera and radar information only leads to a marginal performance improvement of 0.2%. This outcome can be attributed to the nature of the 3D occupancy prediction task, which necessitates dense features. However, the sparse 3D point clouds provided by surround-view radars limit their contribution in rainy scenarios, thus resulting in a negligible impact. In comparison, despite the reflection issue it faces in rainy scenarios, the lidar sensor can still make a substantial contribution to the model’s overall performance owing to its dense 3D point cloud. Another reason could be attributed to the absence of severe rainy conditions in the nuScenes dataset. The dataset primarily includes light to moderate rain scenarios, wherein the lidar data maintains a consistently high quality. By integrating the information from all three sensors, our model achieves the best performance.
In nighttime scenarios, it is undeniable that purely vision-centric approaches perform poorly due to the sensitivity of surround-view cameras to varying illumination conditions. We found that integrating information from surround-view radars notably augmented the model’s performance, leading to an approximate 3% enhancement. Furthermore, the inclusion of radar data significantly improves the capability to predict dynamic objects, and this is even better than merging the lidar data. Particularly, a performance boost of around 10% is observed for the car class. Likewise, we note performance gains of approximately 6% for bicycles and 3% for motorcycles as small dynamic objects. This progress is linked to the velocity measurement function of surround-view radars and demonstrates the effective integration of these features with camera attributes within our framework. What’s more, the inclusion of lidar information yielded an additional performance gain of 6%.
@@ IV-E Challenging Scenarios Performance Analysis
Method Backbone Input Modality Resolution mIoU $\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ const. veh. $\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck $\bullet$ drive. surf. $\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation InverseMatrixVT3D [14] R101-DCN C $200\times 200\times 16$ 27.94 25.08 14.30 43.25 43.32 14.46 21.29 15.93 12.48 28.82 36.40 55.31 23.66 37.75 22.53 18.38 34.06 OccFusion R101+VoxelNet C+R $200\times 200\times 16$ 28.19 27.59 18.73 43.24 48.67 12.05 18.42 15.85 14.46 29.68 38.03 53.62 19.05 35.38 20.89 23.86 31.47 OccFusion R101+VoxelNet C+L $200\times 200\times 16$ 33.24 27.77 17.91 46.36 48.90 22.57 23.43 23.53 22.54 33.75 36.26 49.09 25.53 34.70 23.41 40.58 55.58 OccFusion R101+VoxelNet C+L+R $200\times 200\times 16$ 34.35 29.06 21.53 44.83 54.01 20.00 23.41 23.82 22.42 28.85 42.22 58.41 20.86 38.55 24.57 39.96 57.04
@@
TABLE II: 3D semantic occupancy prediction results on nuScenes validation rainy scenario subset. All methods are trained with dense occupancy labels from [7]. Notion of modality: Camera (C), Lidar (L), Radar (R).

### 09 May 2024 v4

@@ III-C Surround-View Images Feature Extraction
Given surround-view images, we initially employ ResNet101-DCN [54] as our 2D backbone and feature pyramid network (FPN) [55] as the neck to extract multi-scale feature maps. The resulting feature maps have resolutions that are $\frac{1}{8}$, $\frac{1}{16}$, and $\frac{1}{32}$ of the input image resolution, respectively. Subsequently, a view transformation is leveraged, yielding multi-scale global BEV features and local 3D feature volumes. The global and local features with smaller resolutions contain valuable semantic information, aiding the model in predicting the semantic class of each voxel grid. Conversely, those features with larger resolutions provide rich spatial information, enabling the model to determine whether the current voxel grid is occupied or unoccupied.
@@ III-D Lidar Dense 3D Point Cloud Feature Extraction
In this paper, we adopt VoxelNet [32] as our 3D backbone for feature extraction of the 3D point cloud. We begin by voxelizing the 3D point cloud to generate the voxel grid and its associated coordinates. In each voxel grid containing 3D points, 35 points are randomly selected. In each voxel grid containing 3D points, 35 points are randomly selected. If a voxel grid has fewer than 35 points, zero padding is applied to reach 35 points. Each point, denoted as $p_{i}^{Lid}$, includes an initial feature vector $p_{i}^{Lid}=[x_{i},y_{i},z_{i},\gamma_{i}]$ that represents the point’s 3D position and reflectance rate. Subsequently, we calculate the centre position of these 3D points within the same voxel grid and augment each point with a relative offset with respect to the centre position. This augmentation results in a new feature vector $p_{i}^{Lid}=[x_{i},y_{i},z_{i},\gamma_{i},x_{i}-\bar{x},y_{i}-\bar{y},z_{i}-\bar{z}]$. Following this, only the voxel grid with 3D points is input to the 3D backbone to refine the features further, producing the final local 3D feature volume. Additionally, we employ average pooling along the Z-axis of the local 3D feature volume for the global BEV feature to obtain the flattened global BEV feature. The 3D backbone outputs the highest resolution of the global and local features, while the lower resolution features are obtained through 3D/2D downsampling operations.
@@ III-E Radar Sparse 3D Point Cloud Feature Extraction
The Radar 3D point contains richer information compared to the lidar 3D point. Each Radar 3D point has an initial feature vector $p_{i}^{Rad}=[x_{i},y_{i},z_{i},V_{xi},V_{yi}]$, where $V_{xi}$ denote the velocity along the X-axis, and $V_{yi}$ denote the velocity along the Y-axis. Like the lidar point cloud processes, we begin by voxelizing the Radar 3D point cloud and obtaining the voxel grids and their associate coordinates. For the voxel grid with points, we calculate the mean value among the 3D points and augment each point with a relative offset with respect to the mean value. This augmentation results in a new feature vector $p_{i}^{Rad}=[x_{i},y_{i},z_{i},V_{xi},V_{yi},x_{i}-\bar{x},y_{i}-\bar{y},z_{i}-\bar{z},V_{xi}-\bar{V_{x}},V_{yi}-\bar{V_{y}}]$. We subsequently input the non-empty voxel grid into the 3D backbone to obtain the local 3D feature volume and then apply average pooling to obtain the global BEV feature.
@@ III-F Dynamic Fusion 3D/2D
Drawing inspiration from BEVFusion [42, 43] and SENet [56], this research merges two BEV features and two 3D feature volumes by concatenating their feature channels. Subsequently, a Conv3D/2D layer is applied to reduce the feature channel dimension, facilitating the merging of valuable features from diverse modalities while filtering out noisy features. This process is followed by a 3D/2D SENet block, where the merged features are inputted into the squeeze module to determine the importance of each feature channel. The excitation module then implements an excitation procedure by multiplying the merged features with the squeeze feature, enabling critical features to dominate. Details of the Dynamic Fusion 2D module are in the upper section of Figure 3. In contrast, specifics of the Dynamic Fusion 3D module are depicted in the bottom section of Figure 3.
@@ IV-A Implementation Details
The OccFusion utilizes ResNet101-DCN [57, 58] as the 2D backbone, with pre-trained weights provided by FCOS3D [59], to extract image features. The backbone’s feature maps from stages 1, 2, and 3 are then fed into FPN [55], resulting in three levels of multi-scale image features. The network architecture consists of four levels ($L=4$), with no skip connection applied to the highest level. Our framework is adaptable to any view transformation approach. In this paper, we choose the view transformation method proposed in InverseMatrixVT3D [14] to aggregate visual features. It is worth mentioning that when the framework does not merge lidar and radar information, it is the same as the InverseMatrixVT3D algorithm. Thus, OccFusion (C) has the same performance as InverseMatrixVT3D. Our framework uses 10 lidar sweeps, and 5 surround-view radar sweeps for each data sample. To extract 3D feature volumes, the VoxelNet [32] is used as the framework’s 3D backbone, which processes the dense 3D point cloud from lidar and the sparse 3D point cloud from surround-view radars. The AdamW optimizer with an initial learning rate of 5e-5 and weight decay of 0.01 is employed for optimisation. The learning rate is decayed using a multi-step scheduler. The model is trained on eight A10 GPUs, each with 24GB of memory for two days.
IV-B Loss Function
The framework is trained to utilize focal loss [60], Lovasz-Softmax loss [61], and scene-class affinity loss [1]. Considering the significance of high-resolution 3D volumes compared to lower-resolution ones, a decayed loss weight $w=\frac{1}{2^{l}}$ is applied for supervision at the $l$-th level. The ultimate loss formulation is as follows:
$Loss=\sum_{l=0}^{3}\frac{1}{2^{l}}\times(L_{focal}^{l}+L_{lovasz}^{l}+L_{scal\_geo}^{l}+L_{scal\_sem}^{l})$ (2)
where $l$ represent $l$-th level within the framework.
IV-C Dataset
@@
Our 3D semantic occupancy prediction and 3D scene completion experiments were carried out using the nuScenes dataset. The ground truth labels utilized in these experiments were sourced from the works of SurroundOcc [7] and Occ3D [6]. The SurroundOcc’s label spans a range of -50 m to 50 m for the X and Y directions and -5 m to 3 m for the Z direction. This range suits our model’s ablation study on the perception range factor. On the other hand, the Occ3D provide a ground truth label for a relatively minor perceptual range, which spans a range of -40 m to 40 m for the X and Y directions and -1 m to 5.4 m for the Z direction. Furthermore, since Occ3D’s label was designed for purely vision-centric algorithms, a visibility mask is provided for each voxel grid, and evaluation only considers visible voxels. As the test set labels were unavailable, we trained our model on the training set and evaluated its performance on the validation set. Additionally, we chose particular frames from the nuScenes validation set, using ground truth labels from SurroundOcc’s work to establish subsets corresponding to rainy and night scenarios. The distribution of classes in the validation set, rainy scenario subset, and night scenario subset is shown in Figure 4. The class with a sample number equal to zero is not listed in each set.
To further validate the efficacy of our methodology, we conducted a semantic scene completion experiment on the SemanticKITTI dataset utilizing data from the left RGB camera and lidar. SemanticKITTI provides annotated outdoor lidar scans classified into 21 semantic classes. The input image has a resolution of $1241\times 376$, and the ground truth is voxelized into a $256\times 256\times 32$ grid with a voxel size of 0.2m. Given this dataset’s absence of radar sensors, we assessed our OccFusion (C+L) model on the validation set.
IV-D Performance Evaluate Metrics
In assessing the effectiveness of various SOTA algorithms for 3D semantic occupancy prediction and contrasting them with our approach, we utilize Intersection over Union (IoU) to evaluate each semantic class. Moreover, we adopt IoU for occupied voxels, disregarding their semantic class for the scene completion task evaluation. Additionally, the mean IoU across all semantic classes (mIoU) serves as a comprehensive evaluation metric:
@@ IV-D Performance Evaluate Metrics
IV-E Model Performance Analysis
We assess different sensor fusion strategies using our proposed framework on the nuScenes validation set and present the final benchmark results in Table I and Table II. In Table I, compared to purely vision-centric approaches, including sparse 3D point cloud information from the surround-view radars leads to a noteworthy around 2% improvement in performance. Moreover, adding dense 3D point cloud information from the lidar further enhances the performance to approximately 27% mIoU. These experimental results confirm the efficacy of utilizing multi-sensor fusion to substantially improve the performance of the 3D semantic occupancy prediction task. However, in Table II, we observed a non-trivial performance degradation caused by merging radar information. This may contribute to the nature of radar, which is good at measuring objects that are far away, but the perception range in Occ3D’s label and the visibility mask limit its strength.
To further evaluate the effectiveness of our proposed framework, we conduct comparative experiments for the Semantic Scene Completion task on the SemanticKITTI dataset. The benchmark result is presented in Table III. Integrating camera and lidar data in our framework yields highly competitive performance compared to vision-centric and lidar-centric algorithms.
Method Backbone Input Modality IoU mIoU $\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ const. veh. $\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck $\bullet$ drive. surf. $\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation MonoScene [1] R101-DCN C 23.96 7.31 4.03 0.35 8.00 8.04 2.90 0.28 1.16 0.67 4.01 4.35 27.72 5.20 15.13 11.29 9.03 14.86 Atlas [62] R101-DCN C 28.66 15.00 10.64 5.68 19.66 24.94 8.90 8.84 6.47 3.28 10.42 16.21 34.86 15.46 21.89 20.95 11.21 20.54 BEVFormer [3] R101-DCN C 30.50 16.75 14.22 6.58 23.46 28.28 8.66 10.77 6.64 4.05 11.20 17.78 37.28 18.00 22.88 22.17 13.80 22.21 TPVFormer [5] R101-DCN C 30.86 17.10 15.96 5.31 23.86 27.32 9.79 8.74 7.09 5.20 10.97 19.22 38.87 21.25 24.26 23.15 11.73 20.81 C-CONet [63] R101 C 26.10 18.40 18.60 10.00 26.40 27.40 8.60 15.70 13.30 9.70 10.90 20.20 33.00 20.70 21.40 21.80 14.70 21.30 InverseMatrixVT3D [14] R101-DCN C 31.85 18.88 18.39 12.46 26.30 29.11 11.00 15.74 14.78 11.38 13.31 21.61 36.30 19.97 21.26 20.43 11.49 18.47 RenderOcc [9] R101 C 29.20 19.00 19.70 11.20 28.10 28.20 9.80 14.70 11.80 11.90 13.10 20.10 33.20 21.30 22.60 22.30 15.30 20.90 SurroundOcc [7] R101-DCN C 31.49 20.30 20.59 11.68 28.06 30.86 10.70 15.14 14.09 12.06 14.38 22.26 37.29 23.70 24.49 22.77 14.89 21.86 LMSCNet [64] - L 36.60 14.90 13.10 4.50 14.70 22.10 12.60 4.20 7.20 7.10 12.20 11.50 26.30 14.30 21.10 15.20 18.50 34.20 L-CONet [63] - L 39.40 17.70 19.20 4.00 15.10 26.90 6.20 3.80 6.80 6.00 14.10 13.10 39.70 19.10 24.00 23.90 25.10 35.70 M-CONet [63] - C+L 39.20 24.70 24.80 13.00 31.60 34.80 14.60 18.00 20.00 14.70 20.00 26.60 39.20 22.80 26.10 26.00 26.00 37.10 OccFusion R101-DCN C 31.85 18.88 18.39 12.46 26.30 29.11 11.00 15.74 14.78 11.38 13.31 21.61 36.30 19.97 21.26 20.43 11.49 18.47 OccFusion R101+VoxelNet C+R 33.97 20.73 20.46 13.98 27.99 31.52 13.68 18.45 15.79 13.05 13.94 23.84 37.85 19.60 22.41 21.20 16.16 21.81 OccFusion R101+VoxelNet C+L 44.35 26.87 26.67 18.38 32.97 35.81 19.39 22.17 24.48 17.77 21.46 29.67 39.01 21.94 24.90 26.76 28.53 40.03 OccFusion R101+VoxelNet C+L+R 44.66 27.30 27.09 19.56 33.68 36.23 21.66 24.84 25.29 16.33 21.81 30.01 39.53 19.94 24.94 26.45 28.93 40.41
@@ IV-E Model Performance Analysis
Method Backbone Input Modality mIoU $\bullet$ others $\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ const. veh. $\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck $\bullet$ drive. surf. $\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation MonoScene [1] EfficientNetB7 C 6.06 1.75 7.23 4.26 4.93 9.38 5.67 3.98 3.01 5.90 4.45 7.17 14.91 6.32 7.92 7.43 1.01 7.65 BEVDet [2] ResNet101 C 11.73 2.09 15.29 0.0 4.18 12.97 1.35 0.0 0.43 0.13 6.59 6.66 52.72 19.04 26.45 21.78 14.51 15.26 BEVFormer [3] ResNet101 C 23.67 5.03 38.79 9.98 34.41 41.09 13.24 16.50 18.15 17.83 18.66 27.70 48.95 27.73 29.08 25.38 15.41 14.46 BEVStereo [4] ResNet101 C 24.51 5.73 38.41 7.88 38.70 41.20 17.56 17.33 14.69 10.31 16.84 29.62 54.08 28.92 32.68 26.54 18.74 17.49 TPVFormer [5] ResNet101 C 28.34 6.67 39.20 14.24 41.54 46.98 19.21 22.64 17.87 14.54 30.20 35.51 56.18 33.65 35.69 31.61 19.97 16.12 OccFormer [20] ResNet101 C 21.93 5.94 30.29 12.32 34.40 39.17 14.44 16.45 17.22 9.27 13.90 26.36 50.99 30.96 34.66 22.73 6.76 6.97 CTF-Occ [6] ResNet101 C 28.53 8.09 39.33 20.56 38.29 42.24 16.93 24.52 22.72 21.05 22.98 31.11 53.33 33.84 37.98 33.23 20.79 18.00 RenderOcc [9] ResNet101 C 26.11 4.84 31.72 10.72 27.67 26.45 13.87 18.20 17.67 17.84 21.19 23.25 63.20 36.42 46.21 44.26 19.58 20.72 BEVDet4D [10]* Swin-B C 42.02 12.15 49.63 25.10 52.02 54.46 27.87 27.99 28.94 27.23 36.43 42.22 82.31 43.29 54.46 57.90 48.61 43.55 PanoOcc [11]* ResNet101 C 42.13 11.67 50.48 29.64 49.44 55.52 23.29 33.26 30.55 30.99 34.43 42.57 83.31 44.23 54.40 56.04 45.94 40.40 FB-OCC [12]* ResNet101 C 43.41 12.10 50.23 32.31 48.55 52.89 31.20 31.25 30.78 32.33 37.06 40.22 83.34 49.27 57.13 59.88 47.67 41.76 OctreeOcc [13]* ResNet101 C 44.02 11.96 51.70 29.93 53.52 56.77 30.83 33.17 30.65 29.99 37.76 43.87 83.17 44.52 55.45 58.86 49.52 46.33 OccFusion* R101+VoxelNet C+R 38.26 10.11 43.48 26.63 48.25 53.29 24.98 33.35 31.75 27.31 29.88 40.15 77.59 35.47 43.64 48.01 42.73 33.81 OccFusion* R101+VoxelNet C+L 46.79 11.65 47.81 32.07 57.27 57.51 31.80 40.11 47.35 33.74 45.81 50.35 78.79 37.17 44.36 53.36 63.18 63.20 OccFusion* R101+VoxelNet C+L+R 46.67 12.37 50.33 31.53 57.62 58.81 33.97 41.00 47.18 29.67 42.03 48.04 78.39 35.68 47.26 52.74 63.46 63.30
@@
TABLE II: 3D semantic occupancy prediction results on Occ3D benchmark. All methods are trained with dense occupancy labels from [6]. Notion of modality: Camera (C), Lidar (L), Radar (R). ”*” denotes training with the camera mask.
Method IoU mIoU road sidewalk parking other-ground building car truck bicycle motorcycle other-vehicle vegetation trunk terrain person bicyclist motorcyclist fence pole traf.sign LMSCNet [64] 28.61 6.70 40.68 18.22 4.38 0.00 10.31 18.33 0.00 0.00 0.00 0.00 13.66 0.02 20.54 0.00 0.00 0.00 1.21 0.00 0.00 AICNet [65] 29.59 8.31 43.55 20.55 11.97 0.07 12.94 14.71 4.53 0.00 0.00 0.00 15.37 2.90 28.71 0.00 0.00 0.00 2.52 0.06 0.00 3DSketch [66] 33.30 7.50 41.32 21.63 0.00 0.00 14.81 18.59 0.00 0.00 0.00 0.00 19.09 0.00 26.40 0.00 0.00 0.00 0.73 0.00 0.00 JS3C-Net [67] 38.98 10.31 50.49 23.74 11.94 0.07 15.03 24.65 4.41 0.00 0.00 6.15 18.11 4.33 26.86 0.67 0.27 0.00 3.94 3.77 1.45 MonoScene [1] 36.86 11.08 56.52 26.72 14.27 0.46 14.09 23.26 6.98 0.61 0.45 1.48 17.89 2.81 29.64 1.86 1.20 0.00 5.84 4.14 2.25 TPVFormer [5] 35.61 11.36 56.50 25.87 20.60 0.85 13.88 23.81 8.08 0.36 0.05 4.35 16.92 2.26 30.38 0.51 0.89 0.00 5.94 3.14 1.52 InverseMatrixVT3D [14] 36.22 11.81 52.99 25.84 20.04 0.09 13.17 24.08 10.25 1.85 2.65 6.80 16.98 3.09 27.77 4.01 3.13 0.00 4.94 4.05 2.67 VoxFormer [68] 44.02 12.35 54.76 26.35 15.50 0.70 17.65 25.79 5.63 0.59 0.51 3.77 24.39 5.08 29.96 1.78 3.32 0.00 7.64 7.11 4.18 OccFormer [20] 36.50 13.46 58.84 26.88 19.61 0.31 14.40 25.09 25.53 0.81 1.19 8.52 19.63 3.93 32.63 2.78 2.82 0.00 5.61 4.26 2.86 Symphonies [69] 41.44 13.44 55.78 26.77 14.57 0.19 18.76 27.23 15.99 1.44 2.28 9.52 24.50 4.32 28.49 3.19 8.09 0.00 6.18 8.99 5.39 OctreeOcc [13] 44.71 13.12 55.13 26.74 18.68 0.65 18.69 28.07 16.43 0.64 0.71 6.03 25.26 4.89 32.47 2.25 2.57 0.00 4.01 3.72 2.36 UDNet [70] 58.90 20.70 67.00 37.20 20.30 2.20 36.00 42.10 25.70 1.80 2.30 11.20 40.10 18.30 45.80 2.50 1.20 0.00 11.90 23.00 3.80 OccFusion(C+L) 58.68 21.92 65.67 36.33 23.08 0.00 39.09 45.62 20.05 2.96 3.51 8.76 40.68 19.37 45.53 3.16 4.37 0.00 15.70 27.57 15.21
TABLE III: 3D semantic scene completion performance on SemanticKITTI validation set. Notion of modality: Camera (C), Lidar (L)
IV-F Challenging Scenarios Performance Analysis
We assess various sensor fusion strategies in challenging nighttime and rainy scenarios to gain deeper insights into sensor fusion properties and effectiveness. The performance of the models in these scenarios is presented in Table IV and Table V.
In the rainy scenario, despite the sparse 3D point cloud provided by the radar sensors, we observed 2% performance gain by integrating surround-view cameras with radar. Furthermore, despite its reflection issue in rainy scenarios, the lidar sensor can still substantially contribute to the model’s overall performance owing to its dense 3D point cloud. Another reason is the absence of severe rainy conditions in the nuScenes dataset. The dataset primarily includes light to moderate rain scenarios, wherein the lidar data maintains a consistently high quality. Our model performs best by integrating the information from all three sensors.
In nighttime scenarios, it is undeniable that purely vision-centric approaches perform poorly due to the sensitivity of surround-view cameras to varying illumination conditions. We found that integrating information from surround-view radars notably augmented the model’s performance, leading to an approximate 1.2% enhancement. Furthermore, including radar data significantly improves the capability to predict dynamic objects. Particularly, a performance boost of around 4% is observed for the car class. Likewise, we note performance gains of approximately 0.7% for bicycles and 3% for motorcycles as small dynamic objects. This progress is linked to the velocity measurement function of surround-view radars and demonstrates the effective integration of these features with camera attributes within our framework. What’s more, including lidar information yielded an additional performance gain of 4.7%